OpenStudy (anonymous):

PLEASSSSSSE HELPP (x+1)/(x-8) - (x)/(8-x) A) (2x+1)/(x-8) B) (2x-1)/(x-8) C) (2x+1)/(x+8)

OpenStudy (kropot72):

If we multiply the numerator and the denominator of the second expression we get: \[\frac{x+1}{x-8}-\frac{-x}{x-8}\] Now you can use x - 8 as a common denominator.

OpenStudy (kropot72):

\[\frac{x+1-(-x)}{x-8}=you\ can\ finish\]

OpenStudy (anonymous):

thank you so much

OpenStudy (kropot72):

You're welcome :)